OpenStudy (anonymous):

Micah has been following stock ratings. He is keeping his eye on Sam's Pork Bellies, and the pattern can be modeled by a sine function. He has found it to have a maximum of 3, minimum of −5, and period of 2π. Which of the following sine functions could model Micah's stock rating?

OpenStudy (anonymous):

?

OpenStudy (sidsiddhartha):

max=3 and min=-5 ?

OpenStudy (anonymous):

f(x) = 4 sin x − 1 f(x) = 2 sin x − 5 f(x) = 4 sin 2x − 1 f(x) = 3 sin 2x − 5 those are the answers i was thinking d

OpenStudy (anonymous):

is that right?

OpenStudy (sidsiddhartha):

are u sure about the max and min

OpenStudy (anonymous):

im pretty sure.. bc it says it in the sentence

OpenStudy (sidsiddhartha):

yeah ok

OpenStudy (anonymous):

so was i rigt? d?

OpenStudy (anonymous):

i right*

OpenStudy (sidsiddhartha):

yeah so i think f(x) = 4 sin x − 1 because amplitude =avrg(5+3)=4 and let the function f(x)=asin(bx) so 2pi=2pi/b b=1 and a=4

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

sorry im not good with trig...

OpenStudy (anonymous):

like i dont understand if youre saying that 4sin x - 1 is the answer or if youre saying it isnt

OpenStudy (anonymous):

oh wait it is the answer bc the average is the 4!

OpenStudy (anonymous):

right?

OpenStudy (sidsiddhartha):

in the general form f(x)=asin(bx)+c here amplitude, a=4 and b=1andc=-1 that leads f(x)=4sinx-1 thats ur answer

OpenStudy (anonymous):

ok thank you ill post the next one :)

OpenStudy (sidsiddhartha):

ok :)

OpenStudy (anonymous):

Tyyy so muchh I was legit stuck on this one *-*